OpenStudy (anonymous):

Evaluate the following expression using the values given: Find x3 - 2y2 - 3x3 + z4 if x = 3, y = 5, and z = -3

OpenStudy (texaschic101):

just a matter of subbing... x^3 - 2y^2 - 3x^3 + z^4.......x = 3,y = 5, z = -3 3^3 - 2(5^2) - 3(3^3) + -3^4 = can you do the math and finish ?

OpenStudy (anonymous):

so its 27 - 50 - 81 + -81

OpenStudy (texaschic101):

yes...if it makes it easier, write it like this.. 27 - 50 - 81 - 81....do that and you will get you final answer

OpenStudy (anonymous):

so the answer is -185

OpenStudy (texaschic101):

yes

OpenStudy (anonymous):

can i ask another question??

OpenStudy (texaschic101):

sure can :)

OpenStudy (anonymous):

here??

OpenStudy (texaschic101):

yes

OpenStudy (anonymous):

or in another post

OpenStudy (anonymous):

2x2 - 2z4 + y2 - x2 + z4 if x = -4, y = 3, and z = 2

OpenStudy (texaschic101):

same way... 2x^2 - 2z^4 + y^2 - x^2 + z^4....x = -4, y = 3, z = 2 sub 2(-4^2) - 2(2^4) + 3^2 -(-4^2) + 2^4.....you got to be very careful with your signs now, you take it from here....I will stay and check your work. If you have any questions at all, just ask. Take as long as you need.

OpenStudy (texaschic101):

just remember, a neg times a neg is a positive

OpenStudy (texaschic101):

you need help ? You need me to start you off ?

OpenStudy (anonymous):

-16 -16 + 9 - 8 + 16

OpenStudy (texaschic101):

2(16) - 2(16) + 9 - 16 + 16 ....you following this ??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so the answer is 9

OpenStudy (texaschic101):

now there is several things that cancel out here.....like the 2(16) and the -2(16) and the -16 and the + 16......leaving you with just 9 as your answer.

OpenStudy (texaschic101):

yes..it is 9

OpenStudy (anonymous):

thank you very much

OpenStudy (texaschic101):

glad to help :)